### v1

@@ 5 Environmental Requirements
It is desirable to have a maintenance-free operation and a longer service time. Safety is an important feature that should be tested under different operational conditions, as discussed in the next section. Hence, to achieve these goals, well-regulated standards have been designed and are in place as environmental standards. The idea was to test the components under different environmental conditions to verify the effectiveness of the product. Mechanical shock and vibration, heat and humidity, chemical contamination, and other tests were conducted. The tests are performed at the vehicle and subsystem levels. Usually, automakers are responsible for vehicle-level testing and suppliers are responsible for component-level testing. At the vehicle level, a widely accepted regulation has been developed under the United Nations (UN), which is called ECE 100. A vehicle with an electric powertrain should comply with this regulation. In addition, the battery pack was covered by this standardECE100R2 . This standard also covers safety concerns. At the component level, battery packs or parts such as the BMS and ISO 16750 series (road vehicles, – environmental conditions, and testing for electrical and electronic equipment) are very common ISO16750-1 ; ISO16750-2 ; ISO16750-3 ; ISO16750-4 ; ISO16750-5 . ISO 16750 was originally developed for engine-based vehicles but is still used for hybrid and electric vehicles. More widely accepted EV-dedicated standards are expected shortly. The second part of ISO 16750 considers the electrical specifications and is called the electrical load in the standard ISO16750-2 . For example, a BMS board can be the subject of this test. The power supply variation, reverse polarity connection of the device, open-circuit tests, short-circuit protection, and insulation resistance are part of this section. The third part of the standard is devoted to the mechanical aspects ISO16750-3 . Vibration, mechanical shock, and stress are the subjects of this standard. Part four of the standard is called climatic loads, which concern temperature, humidity, dust, corrosion, salt water, ice tests, and similar topics. The most difficult part of the test is the combined humidity and vibration test, which is very challenging. Depending on the installation location inside the vehicle, different maximum values were tested. For example, the climate temperature range is $-45^{0}C$ to $50^{o}C$ but if a component is close to the internal combustion engine, the maximum temperature will be increased to $80^{o}C$ for testing. Another case is when the vehicle is under direct sunlight for a while, and the temperature is increased more than in the environment. Hence, the standard for the value of temperature and cycle testing considers the class of the component that should be referred to. The final part of the ISO 16750 standard is dedicated to chemical issues ISO16750-5 . Electronic boards or devices are subject to different chemical materials for a predetermined time or cycle. The concept is that the component should perform as stated in the standard during or after the test. In most cases, a visual check is required to detect possible damage to a component. As mentioned in this section, there are different types of environmental tests to ensure the performance of the component for different cases. For example, the pack drop test, water immersion, and similar tests are part of the ISO 1670 series or comparable standards. In addition, other standards regulate other aspects, such as production or labeling requirements, or do not use toxic materials. For example, one can refer to lead-free soldering materials or bans on cadmium materials in battery cells.
6 Safety
@@
Battery safety is an important aspect of EVs, as an emerging technology. There have been several reports of EV accidents Crash . There are different safety layers and standard regulations from the vehicle level to the sub-component level. In response to fulfilling standard regulations and achieving a safe system, the ISO 25265 series has been in place, followed by automakers and suppliers ISO26262-1 ; ISO26262-2 ; ISO26262-3 ; ISO26262-4 ; ISO26262-5 ; ISO26262-6 ; ISO26262-7 ; ISO26262-8 ; ISO26262-9 ; ISO26262-10 ; ISO26262-11 ; ISO26262-12 . The topic is briefly reviewed in this section, and it is expected that standards or relevant references will be visited for more information.
@@ 7.1 Cell Technology
Li-ions are the dominant chemical species in automotive battery cells. There are different types of chemistries, such as Lithium Iron Phosphate (LFP), Lithium Nickel Manganese Cobalt Oxide (NMC), and Lithium Cobalt Oxide (LCO), to name but a few. However, there are more details in the chemistry that are not publicly available. The NMC and LFP cells are widely used in automotive applications. The LFP has a lower voltage but better temperature stability. Battery cells are mainly available in three packages: pouch, cylindrical, and prismatic. The pouch is easier to mass-produce and cool and is used by most automakers. In contrast, Tesla uses cylinder cells that require a more complicated cooling structure and better safety. As mentioned earlier, cell chemistries are very different, and it is difficult to classify them accurately. Examples of main cell manufacturers for automotive applications are LG, CATL, and Panasonic. More companies and new companies, such as Northvolt, will emerge as well. The ideal cell should be sufficiently high to provide the required ampere. For the above-mentioned 100 kWh pack, the ideal single-cell capacity is 250 Ah. However, the available cell in the automotive is approximately 80 Ah for the pouch type. Hence, three cells were be connected in parallel to provide the required $Ah$. Pouch cells with a capacity of approximately 200 Ah are under development, and it is expected to see them in a couple of years.
@@ 7.3 Battery Electric Power Architecture
8 BMS
The BMS in the battery pack has several functions, including monitoring of cells, monitoring of the temperature at some points, controlling the contactors, monitoring of the pack current, communication towards the vehicle, and additional functions that will be discussed in this section. The idea was to complement the materials provided in the previous sections. BMS play a critical role in ensuring the optimal performance, safety, and longevity of electric vehicle battery packs. Among the primary tasks of a BMS are the monitoring and estimation of the State of Charge (SOC), State of Health (SOH), and temperature of the battery cells. These parameters are crucial for maintaining efficient and reliable operation.
@@ 8.1 Module Monitoring and Control Unit
Fig. 6 shows the BMS system, including a daisy communication channel between the modules and the controller. Each module had a monitoring board in which the voltage of the cells and a couple of temperature points of the modules were constantly measured. As mentioned earlier, there are parallel cell connections in the first place. Subsequently, the common voltage was measured. The individual current distribution of cells is not measured owing to the complexity and cost impact. The estimation of the state of charge or available power of parallel-connected cells requires more work, as shown in an example in Han . Temperature measurement was performed inside the module at appropriate points. It can be a couple of points inside the module or more points, including all cells. More points add more complexity, especially when considering the ASIL level of the signals. This adds more limitations to the system. A practical compromise may be the measurement of two points between the cells to ensure that the system does not face a thermal runaway. Another feature of the module controller unit is cell balancing. There are passive and active cell balancing methods Naguib ; Ghaeminezhad , in which passive cell balancing is common for passenger cars because of its simplicity and cost. Because of the large number of cells in the pack, collecting and processing the required information is a challenging task. The module controller collects all measurements and sends them to the electronic control unit (ECU) via a serial bus, such as isoSPI. In this manner, each module has a pair of wires as the input and another pair of wires as the output. As shown in Fig. 6, a serial loop was formed, including all module controllers and the ECU. An ECU may have two connection sides. Usually, an ASIL D level is required for communication, and if one cable is disconnected, the second cable can continue the operation. Hence, in this ring configuration, it is possible to reach each node on two paths. One important development is the removal of all wiring between the cells and communication and the use of a wireless structure. This activity is ongoing and it is expected to see more products in the market in the near future. One major issue with a battery pack is the integrity of the communication link.
8.2 ECU
@@ 9 Other Topics and New Trends
The specific energy and power are important parameters in the cell and pack levels that can be countered as performance indices. As mentioned in the previous section, the first step is to construct battery modules as independent units and then integrate them into a pack to fulfill all requirements and standard regulations. The general trend is towards more compact and lighter systems with reduced costs. For pouch cells, an energy density of 650 Wh/L and specific energy of 270 Wh/kg can be found in vehicles. For upcoming cells, a 10-20% enhancement will appear in a couple of years. As an approximation, these numbers can be reduced by approximately 40% at the pack level. For example, an automotive battery pack using cells with 270 Wh/kg, can have an energy density of approximately 180 Wh/kg. New cell technologies with improved energy density or specific energy numbers, more integration in packaging, and improved technology for other parts of the system, such as contactors, sensors, and elimination of wires (using wireless), are needed to reduce the weight and volume of the pack. Hence, new materials and chemistry play key roles in meaningful changes in technological advancement. The battery system voltage is another interesting topic. A nominal voltage of 350 V is widely used for passenger cars. However, in some cases, the system voltage was 700-800 V. The power requirement has increased from less than a hundred kW to more than 200 kW and in some cases, it has increased even further. Consequently, it is necessary to increase the system voltage level to reduce the current level in the system. Consequently, several high-voltage developments such as 800 V is under development by the automotive industry, which will increasingly appear in the market Haghbin2 . Another point that should be mentioned is the charging infrastructure. Increasing the battery voltage may add issues to the available fast-charging infrastructure. Some fast chargers can not charge 800 V batteries. Hence, there are some works under development to change the classical structure such that the vehicle can utilize the low-voltage charger to charge high-voltage packs. In addition to passenger cars, battery packs are required for buses, trucks, and aviation vehicles. Usually, the solutions developed for passenger cars are optimized, and the results are expected to be utilized in other sections. For aviation electrification, some major hindrances require special attention and further development. The classical structure of a battery pack is inappropriate for aviation. It is not possible to disconnect the battery pack in the case of an extreme fault, and new architectures are required. In addition, the environmental conditions are harsher and packaging may be more challenging. Available aviation solutions, such as satellites, are extremely expensive and not affordable. Hence, more work and maturity are required in this sector.

### v2

@@ 5 Environmental Requirements
It is desirable to have a maintenance-free operation and a longer service time. Safety is an important feature that should be tested under different operational conditions, as discussed in the next section. Hence, to achieve these goals, well-regulated standards have been designed and are in place as environmental standards. The idea was to test the components under different environmental conditions to verify the effectiveness of the product. Mechanical shock and vibration, heat and humidity, chemical contamination, and other tests were conducted. The tests are performed at the vehicle and subsystem levels. Usually, automakers are responsible for vehicle-level testing and suppliers are responsible for component-level testing. At the vehicle level, a widely accepted regulation has been developed under the United Nations (UN), which is called ECE 100. A vehicle with an electric powertrain should comply with this regulation. In addition, the battery pack was covered by this standardECE100R2 . This standard also covers safety concerns. At the component level, battery packs or parts such as the BMS and ISO 16750 series (road vehicles, – environmental conditions, and testing for electrical and electronic equipment) are very common ISO16750-1 ; ISO16750-2 ; ISO16750-3 ; ISO16750-4 ; ISO16750-5 . ISO 16750 was originally developed for engine-based vehicles but is still used for hybrid and EVs. More widely accepted EV-dedicated standards are expected shortly. The second part of ISO 16750 considers the electrical specifications and is called the electrical load in the standard ISO16750-2 . For example, a BMS board can be the subject of this test. The power supply variation, reverse polarity connection of the device, open-circuit tests, short-circuit protection, and insulation resistance are part of this section. The third part of the standard is devoted to the mechanical aspects ISO16750-3 . Vibration, mechanical shock, and stress are the subjects of this standard. Part four of the standard is called climatic loads, which concern temperature, humidity, dust, corrosion, salt water, ice tests, and similar topics. The most difficult part of the test is the combined humidity and vibration test, which is very challenging. Depending on the installation location inside the vehicle, different maximum values were tested. For example, the climate temperature range is $-45^{0}C$ to $50^{o}C$ but if a component is close to the internal combustion engine, the maximum temperature will be increased to $80^{o}C$ for testing. Another case is when the vehicle is under direct sunlight for a while, and the temperature is increased more than in the environment. Hence, the standard for the value of temperature and cycle testing considers the class of the component that should be referred to. The final part of the ISO 16750 standard is dedicated to chemical issues ISO16750-5 . Electronic boards or devices are subject to different chemical materials for a predetermined time or cycle. The concept is that the component should perform as stated in the standard during or after the test. In most cases, a visual check is required to detect possible damage to a component. As mentioned in this section, there are different types of environmental tests to ensure the performance of the component for different cases. For example, the pack drop test, water immersion, and similar tests are part of the ISO 1670 series or comparable standards. In addition, other standards regulate other aspects, such as production or labeling requirements, or do not use toxic materials. For example, one can refer to lead-free soldering materials or bans on cadmium materials in battery cells.
6 Battery Thermal Management system (BTMS) and Safety
Battery thermal management and safety is an important aspect of EVs, as an emerging technology. One of the major safety concerns is thermal runaway, which can occur due to excessive heat, internal short circuits, overcharging, or mechanical damage to the battery cells. This can lead to battery pack failure, fire, or even explosion, posing a significant hazard to the vehicle, passengers, and the surrounding environment.
To mitigate the risks associated with thermal runaway, Battery Thermal Management Systems (BTMS) are employed. These systems are designed to maintain safe operating temperatures across the battery pack, preventing overheating and minimizing the likelihood of thermal runaway events. Proper temperature regulation is essential for ensuring the performance, longevity, and safety of battery packs. There are various technologies used within BTMS to control the temperature of the battery pack R3_Comprehesive :
  • Active and Passive Cooling: Active cooling systems use pumps and liquid coolants to remove excess heat from the battery cells, ensuring that the cells remain within their safe temperature range. Passive cooling systems, on the other hand, use natural convection, heat pipes, or other methods to dissipate heat without the use of pumps or external power sources. Both active and passive cooling mechanisms play a vital role in preventing overheating of the battery pack. These systems ensure that the battery operates within safe thermal limits during high-demand scenarios, such as fast charging or high-power discharges, and thus help prevent the initiation of thermal runaway R3_thermal ; R3_Comprehesive ; R3_development .
  • Phase-Change Materials (PCMs): PCMs are substances that absorb and store heat as they change phase (from solid to liquid) at a specific temperature. In a BTMS, PCMs are integrated into the battery pack to help absorb excess heat during periods of high charge or discharge rates. This helps maintain a more uniform temperature across the battery pack and prevents localized overheating. As the temperature rises, the PCM absorbs thermal energy and transitions to a liquid state, stabilizing the pack’s temperature. When the temperature falls, the material solidifies, releasing stored heat. This capability can significantly enhance the safety and efficiency of EV battery packs R3_development .
  • Immersion Cooling: Immersion cooling involves submerging the battery pack in a thermally conductive liquid that directly absorbs heat from the cells. This method of thermal management provides more uniform and efficient cooling compared to traditional air cooling systems. Immersion cooling is especially useful for high-power applications, where rapid temperature fluctuations are more likely. By submerging the cells in a liquid coolant, the heat is more effectively transferred from the cells to the coolant, reducing the chance of localized overheating and thermal runaway R3_thermal .
@@ 6 Battery Thermal Management system (BTMS) and Safety
In addition to these technologies, BMS also play a key role in ensuring safety. They continuously monitor the battery’s voltage, temperature, and other critical parameters, triggering corrective actions when abnormal conditions are detected. For example, the BMS can activate the cooling system or limit charging rates to prevent overheating. There have been several reports of EV accidents Crash . There are different safety layers and standard regulations from the vehicle level to the sub-component level. In response to fulfilling standard regulations and achieving a safe system, the ISO 25265 series has been in place, followed by automakers and suppliers ISO26262-1 ; ISO26262-2 ; ISO26262-3 ; ISO26262-4 ; ISO26262-5 ; ISO26262-6 ; ISO26262-7 ; ISO26262-8 ; ISO26262-9 ; ISO26262-10 ; ISO26262-11 ; ISO26262-12 . The topic is briefly reviewed in this section, and it is expected that standards or relevant references will be visited for more information.
@@ 7.1 Cell Technology
Li-ions are the dominant chemical species in automotive battery cells. There are different types of chemistries, such as Lithium Cobalt Oxide (LCO), Lithium Iron Phosphate (LFP), Lithium Nickel Cobalt Aluminum Oxide (NCA), and Lithium Nickel Manganese Cobalt Oxide (NMC), to name but a few. However, there are more details in the chemistry that are not publicly available. Fig. 5 shows LCO, LFP, NCA, and NMC battery types, which are different in the cathode types and have different nominal voltage, and so on. The charactristics of these batteries are presented in Table 2.
Figure 5: Battery cells based on Cathode types. (a) LCO. (b) LFP. (c) NCA. (d) NMC Barkholtz .
Table 2: Li-ion battery technologies characteristics Barkholtz ; Larijani_book .
Battery LCO LFP NCA NMC
Cathode $\mbox{LiCoO}_{2}$ $\mbox{LiFePO}_{4}$ $\mbox{LiNi}_{0.8}\mbox{Co}_{0.15}\mbox{Al}_{0.05}\mbox{O}_{2}$ $\mbox{LiNi}_{0.8}\mbox{Mn}_{0.15}\mbox{Co}_{0.05}\mbox{O}_{2}$
Nominal capacity ($Ah$) 2.5 1.1 2.9 3
Nominal voltage ($V$) 2.5 1.1 3.6 3.6
Maximum discharge 20 30 6 20
current ($A$)
Mass ($g$) 44.9 39.4 44.8 44.9
Volumetric energy density 533.4 212.1 676 612.1
($Wh/L$)
Specific energy density 195.8 88.9 243 224.8
($Wh/kg$)
Operating temperature 0 to 50 -20 to 55 0 to 45 0 to 50
range under charge (${}^{\circ}C$)
Operating temperature -20 to 55 -30 to 55 -20 to 60 -20 to 55
range for discharge (${}^{\circ}C$)
NMC and LFP cells are widely used in automotive applications. The LFP has a lower voltage but better temperature stability. Battery cells are available mainly in three packages: pouch, cylindrical, and prismatic, as shown in Fig. 6. The pouch is easier to produce and to cool in mass and is used by most automakers. In contrast, Tesla uses cylinder cells that require a more complicated cooling structure and better safety. As mentioned earlier, cell chemistries are very different and it is difficult to classify them accurately. Examples of main cell manufacturers for automotive applications are LG, CATL, and Panasonic. More companies and new companies, such as Northvolt, will emerge as well. The ideal cell should be sufficiently high to provide the required ampere. For the above-mentioned 100 kWh pack, the ideal single-cell capacity is 250 Ah. However, the available cell in the automotive is approximately 80 Ah for the pouch type. Hence, three cells were be connected in parallel to provide the required $Ah$. Pouch cells with a capacity of approximately 200 Ah are under development, and it is expected to see them in a couple of years.
Figure 6: Battery cell packages. (a) cylindrical. (b) pouch. (c) prismatic.
@@ 7.3 Battery Electric Power Architecture
8 Battery Management System (BMS)
The BMS in the battery pack has several functions, including monitoring of cells and the temperature at some points, controlling the contactors, monitoring of the pack current, communication towards the vehicle, and additional functions that will be discussed in this section. The idea was to complement the materials provided in the previous sections. BMS play a critical role in ensuring the optimal performance, safety, and longevity of EV battery packs. Among the primary tasks of the BMS are the SOC estimation, which determines the remaining charge in the battery; SOH estimation, which evaluates the overall health of the battery; and temperature monitoring, which ensures that the battery operates within safe temperature limits. These parameters are crucial for maintaining efficient and reliable operation.
8.1 Estimation methods
8.1.1 State-of-Charge (SoC) Estimation
The SOC estimation is crucial for ensuring that the battery is neither overcharged nor excessively discharged, both of which can shorten its lifespan. The most common methods for SOC estimation include the following:
  1. 1.
    Coulomb-counting Method: This classical method, which can also be called the ampere count method, calculates the SoC by integrating the current over time Larijani . Although simple and computationally efficient, it can suffer from errors due to the accumulation of drift over time, requiring periodic recalibration or corrections. In addition, the initial SoC should be known to accurately compute the SoC over time Edorado_2 .
  2. 2.
    Model-Based Methods: Recent advancements have introduced more sophisticated methods, such as those based on Kalman filters, which dynamically adjust to changing conditions. One notable study by Locorotondo et al. Locorotondo demonstrated the use of model-adaptive Kalman filters to estimate SoC more accurately by adjusting the model parameters based on real-time data, such as current and voltage measurements. These methods can reduce the error margins seen in simpler techniques and provide more reliable estimates of SoC.
  3. 3.
    Spectroscopy Methods: These methods use techniques such as impedance spectroscopy to estimate SoC by analyzing the battery’s internal resistance and other electrochemical properties spect . While these methods can provide high precision, they tend to be more complex and require specialized equipment.
8.1.2 State of Health (SoH) Estimation
The SoH refers to the ability of the battery to perform at its rated capacity and maintain charge retention over time. Model-based approaches, similar to those used for SoC estimation, are commonly employed for SoH estimation. These models often incorporate aging effects and degradation patterns, providing more accurate forecasts of battery performance over their lifecycles battery_degradation ,Edorado . The temperature is a critical parameter for maintaining battery safety and performance. Excessive temperatures can accelerate aging, whereas low temperatures can reduce efficiency. The BMS typically include temperature sensors located at multiple points in the battery pack to monitor these fluctuations. Advanced methods for temperature management often involve combining sensor data with predictive models to optimize the battery’s thermal management system in real-time. Despite significant progress, several challenges remain in the accurate estimation of SoC, SoH, and temperature. For example, temperature variations across cells and the aging effects of battery components can complicate estimation processes. Recent advancements in machine learning and adaptive algorithms show promise in improving the accuracy of these estimations. By improving the methods for estimating these parameters, modern BMS contribute significantly to the safety, efficiency, and durability of EV batteries, ensuring that they operate within optimal conditions throughout their lifecycle.
@@ 8.2 Module Monitoring and Control Unit
Fig. 8 shows the BMS, including a daisy communication channel between the modules and the controller. Each module had a monitoring board in which the voltage of the cells and a couple of temperature points of the modules were constantly measured. As mentioned earlier, there are parallel cell connections in the first place. Subsequently, the common voltage was measured. The individual current distribution of cells is not measured owing to the complexity and cost impact. The estimation of the state of charge or available power of parallel-connected cells requires more work, as shown in an example in Han . Temperature measurement was performed inside the module at appropriate points. It can be a couple of points inside the module or more points, including all cells. More points add more complexity, especially when considering the ASIL level of the signals. This adds more limitations to the system. A practical compromise may be the measurement of two points between the cells to ensure that the system does not face a thermal runaway. Another feature of the module controller unit is cell balancing. There are passive and active cell balancing methods Naguib ; Ghaeminezhad , in which passive cell balancing is common for passenger cars because of its simplicity and cost. Because of the large number of cells in the pack, collecting and processing the required information is a challenging task. The module controller collects all measurements and sends them to the electronic control unit (ECU) via a serial bus, such as isoSPI. In this manner, each module has a pair of wires as the input and another pair of wires as the output. As shown in Fig. 8, a serial loop was formed, including all module controllers and the ECU. An ECU may have two connection sides. Usually, an ASIL D level is required for communication, and if one cable is disconnected, the second cable can continue the operation. Hence, in this ring configuration, it is possible to reach each node on two paths. One important development is the removal of all wiring between the cells and communication and the use of a wireless structure. This activity is ongoing and it is expected to see more products in the market in the near future. One major issue with a battery pack is the integrity of the communication link.
8.3 Electronic Control Unit (ECU)
@@ 9 Other Topics and New Trends
The specific energy and power are important parameters in the cell and pack levels that can be countered as performance indices. As mentioned in the previous section, the first step is to construct battery modules as independent units and then integrate them into a pack to fulfill all requirements and standard regulations. The general trend is towards more compact and lighter systems with reduced costs. For pouch cells, an energy density of 650 Wh/L and specific energy of 270 Wh/kg can be found in vehicles. For upcoming cells, a 10-20% enhancement will appear in a couple of years. As an approximation, these numbers can be reduced by approximately 40% at the pack level. For example, an automotive battery pack using cells with 270 Wh/kg, can have an energy density of approximately 180 Wh/kg. New cell technologies with improved energy density or specific energy numbers, more integration in packaging, and improved technology for other parts of the system, such as contactors, sensors, and elimination of wires (using wireless), are needed to reduce the weight and volume of the pack. Hence, new materials and chemistry play key roles in meaningful changes in technological advancement. The battery system voltage is another interesting topic. A nominal voltage of 350 V is widely used for passenger cars. However, in some cases, the system voltage was 700-800 V. The power requirement has increased from less than a hundred kW to more than 200 kW and in some cases, it has increased even further. Consequently, it is necessary to increase the system voltage level to reduce the current level in the system. Consequently, several high-voltage developments such as 800 V is under development by the automotive industry, which will increasingly appear in the market Haghbin2 . Another point that should be mentioned is the charging infrastructure.
Another interesting area of research and development in the battery sector involves sodium-based cells, which present an alternative to lithium-ion batteries. Sodium-based batteries utilize sodium ions instead of lithium ions, offering potential advantages such as abundant raw material availability and lower costs, reducing the environmental and ethical concerns associated with lithium and cobalt mining. These sodium-ion batteries are gaining attention for their potential to power EVs in the future, especially as the demand for sustainable and cost-effective solutions increases. Although still in the early stages compared to lithium-ion batteries, sodium-based cells are expected to play a role in the evolution of EV batteries TNO . Another emerging trend is the development of batteries that do not contain rare earth elements. Many current EV battery chemistries, such as those using lithium cobalt oxide, rely heavily on rare earth materials. However, as the industry moves toward more sustainable practices, researchers are focusing on new technologies that minimize or eliminate the use of rare earths. These technologies aim to reduce environmental impact and alleviate concerns about supply chain limitations and geopolitical issues surrounding the extraction of these critical materials. TNO battery technology, which is characterized by high charging and discharging power capabilities, is also gaining attention. This innovation provides rapid charge/discharge cycles, offering the potential for reducing charging times significantly and improving vehicle performance. As the automotive industry continues to focus on reducing charging times and enhancing power delivery, TNO’s battery technology could become a key player in meeting these demands. Additionally, battery swapping systems, as introduced by companies like NIO, offer a unique solution to the issue of long charging times NIO . Instead of waiting for a battery to charge, NIO offers the ability to swap out the battery entirely in a matter of minutes. This system addresses the challenge of limited fast-charging infrastructure and provides a flexible solution for EV owners, particularly in regions with growing EV adoption but underdeveloped charging networks. Increasing the battery voltage may add issues to the available fast-charging infrastructure. Some fast chargers can not charge 800 V batteries. Hence, there are some works under development to change the classical structure such that the vehicle can utilize the low-voltage charger to charge high-voltage packs. In addition to passenger cars, battery packs are required for buses, trucks, and aviation vehicles. Usually, the solutions developed for passenger cars are optimized, and the results are expected to be utilized in other sections. For aviation electrification, some major hindrances require special attention and further development. The classical structure of a battery pack is inappropriate for aviation. It is not possible to disconnect the battery pack in the case of an extreme fault, and new architectures are required. In addition, the environmental conditions are harsher and packaging may be more challenging. Available aviation solutions, such as satellites, are extremely expensive and not affordable. Hence, more work and maturity are required in this sector.